OpenStudy (anonymous):

Looking for help with solving an ODE with Laplace Transform. I know what I'm doing and I have the solution but I think a huge step was skipped and I don't understand how they got what they got.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you familiar with ?

OpenStudy (kinggeorge):

Could you post the solution you have, or at least the part where you think a step was skipped?

OpenStudy (anonymous):

Or the question itself?

OpenStudy (anonymous):

The DE: \[EI \frac{ d^{4}y }{ dx^{4} } = W_{0} \delta(x-L/2);\] \[y(0) = 0; y'(0) = 0; y''(L) = 0; y'''(L) = 0\] The solution: http://img694.imageshack.us/img694/848/solution13m.jpg I don't know how to use y''(L) and y'''(L)

OpenStudy (anonymous):

number 13

OpenStudy (kinggeorge):

What is \(P_0\)?

OpenStudy (anonymous):

Wo

OpenStudy (anonymous):

different textbook editions lol

OpenStudy (kinggeorge):

And the \(\delta\) is the dirac delta?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i just don't know how to use y''(L) and y'''(L) to get what he got

OpenStudy (kinggeorge):

I think, they just integrated 0 with respect to s. So \[\int y^{(4)} ds=y^{(3)}+C\]And since \(y^{(4)}=0\) except at L/2, and \(y^{(3)}(L)=0\), we have that \(y^{(3)}(0)=W_0/EI\). Or something like that. And if you do it again, you get that \(y''(0)=W_0L/2EI\).

OpenStudy (kinggeorge):

This seems a bit hand-wavy to me, but I think they've done something like that.

OpenStudy (anonymous):

Why is \[y^{(4)} = 0\] except at L/2 ?

OpenStudy (anonymous):

oh. because of the delta dirac function

OpenStudy (anonymous):

at L/2 its inf

OpenStudy (kinggeorge):

Basically.

OpenStudy (kinggeorge):

And since you're multiplying by \(W_0/EI\), when you integrate, you get \(W_0/EI\) I think. I'm still not sure why you have to divide by 2 when you do it again.

OpenStudy (anonymous):

thanks for the help :) king indeed

OpenStudy (kinggeorge):

You're welcome.